Intregation of x7dx/(x12 - 1).
Please solve it as soon as possible.

Grade:11

1 Answers

Aditya Gupta
4 years ago
simply substitute x^4=y so that 4x^3dx= dy. also y^3= x^12
now x7dx/(x12 - 1).= x4*(x3dx)/(x12 - 1).= ydy/4(y^3 – 1)
now write the den y^3 – 1 as (y-1)(y^2+y+1).
and express y/(y-1)(y^2+y+1) in partial fractions as A/(y-1) + (By+C)/(y^2+y+1).
after that it is in standard forms and can be easily solved.
kindly approve :)